OpenStudy (anonymous):

250cm^3 OF 0.2 MOLAR POTASSIUM SULPHATE SOLUTION IS MIXED WITH25OCM^3 OF 0.2 MOLAR KCL SOLUTION.THE MOLAR CONCENTRATION OF K POSITIVE IONS IS 0.2 MOLAR 0.25 MOLAR 0.3 MOLAR O.35 MOLAR

OpenStudy (anonymous):

First we calculate the moles of Potassium Sulfate(Sulphate). \[250cm^3=0.25L\] To calculate moles: \[0.25L*0.2\frac{mol}{L}=0.05mol K_2SO_4\] Now that we have the moles of K_2SO_4 we can determine the moles of K. \[0.05molK_2SO_4\frac{2molK}{1molK_2SO_4}=0.1molK\] Time to start working on the second solution: \[0.25L*0.2\frac{mol}{L}=0.05mol KCl\] Total moles of K in our final solution: \[0.1molK+0.05molK=0.15molK\] Our total solution is: \[0.25L+0.25L=0.5L\] From this we can calculate the Molarity: \[\frac{mol}{V}=M\] \[\frac{0.15mol}{0.5}=M\] \[M=0.3\frac{mol}{L}=0.3M\]